# How Similar Are Grokipedia and Wikipedia? A Multi-Dimensional Textual and Structural Comparison

### v3

@@ 3.6 Comparison within Topics
Because these articles cover a diverse set of subject areas, we additionally classified each article using GPT-5 (see Methods for details). This allows us to compare similarity and bias shift not only across individual articles but also across broader thematic domains.
@@
Figure 8: Topic-level comparison of similarity and political bias shift between Wikipedia and Grokipedia articles. (a) Mean combined similarity scores between Wikipedia and Grokipedia articles within each topic, with standard-error bars. (b) Mean political bias shift (Grokipedia – Wikipedia) for articles in each topic, indicating whether Grokipedia content tends to lean more left or right relative to Wikipedia. Error bars show standard errors within each topic.
@@ 4 Discussion and Conclusion
This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across roughly 17,700 matched article pairs drawn from the 20,000 most-edited English Wikipedia titles. Despite their distinct production paradigms, community-driven versus AI-generated, the two encyclopedias exhibit remarkable alignment at the level of meaning and style. Semantic similarity averages around 0.86 and stylistic similarity around 0.88, indicating that Grokipedia reproduces much of Wikipedia’s linguistic and conceptual structure. However, the platforms diverge substantially in form and informational scaffolding. Grokipedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels and lexical diversity (greater syntactic complexity), but reference density. This pattern suggests that Grokipedia’s generation process elaborates on existing material, expanding text length and rhetorical flow, rather than producing substantively new or more rigorously sourced knowledge.
@@
The similarity distributions reveal additional nuance. The bimodal pattern across measures indicates that while some Grokipedia pages closely mirror Wikipedia’s organisation and tone, others diverge sharply. This heterogeneity suggests a hybrid authorship logic: part replication, part improvisation. In some cases, Grok appears to reproduce Wikipedia’s structure with high fidelity; in others, it departs substantially from the source material, generating content that is stylistically and semantically distinct.
@@
Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [apnews2025, businessinsider2025, yasseri2025conversation]. Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia, often copied or paraphrased, while selectively emphasizing Musk’s personal and political narratives [rogers2025wired, oremus2025, hartwelle2025, wong2025]. Our quantitative results support this view: Grokipedia achieves high semantic alignment with Wikipedia but introduces stylistic inflation and reduced citation density. In other words, Grokipedia’s output largely repackages existing human-curated content through an AI lens that privileges fluency and narrative over attribution. Rather than “purging bias,” the system appears to re-encode it, subtly and opaquely, within model parameters and prompt conditioning.
@@
Complementing these textual and stylistic findings, our reference-level analysis (Figures 65) reveals that the magnitude and direction of political bias shifts between Grokipedia and Wikipedia articles are strongly related to article similarity. When Grokipedia reproduces Wikipedia closely (high similarity), the resulting entries show little or no bias shift. However, low-similarity articles—those most substantially rewritten—display both greater variance and a systematic rightward tendency in bias shift. This pattern suggests that editorial divergence in Grokipedia is concentrated precisely where generative rewriting occurs.
The corresponding breakdown of reference sources further supports this interpretation. In the low-similarity subset, citation networks are dominated by a narrower set of news outlets with identifiable political leanings, as measured by the News Media Bias and Factuality dataset [sanchez2024mapping]. Whereas Wikipedia’s reference base remains diverse and balanced, Grokipedia’s alternative sourcing emphasizes ideologically marked or partisan media domains. Together, these results indicate that Grokipedia’s generative process may amplify latent source bias when departing from human-authored baselines, substituting breadth and attribution with ideologically skewed curation.
@@
Several limitations qualify these findings. First, the dataset focuses on Wikipedia’s 20,000 most-edited pages, of which around 17,700 have counterparts on Grokipedia. This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [yasseri2012dynamics]. Second, the similarity metrics employed, lexical, semantic, structural, and stylistic, quantify form and textual alignment but not factual accuracy or ideological framing. The presence of hallucinated claims, selective omissions, or subtle rhetorical shifts remains outside the scope of automated comparison. Third, both platforms are dynamic: Wikipedia continuously evolves, while Grokipedia’s generative parameters may change with future model retraining. Finally, Grokipedia’s underlying data sources and editorial interventions are opaque, preventing full provenance auditing or causal inference about model bias.
@@
These findings extend beyond a single platform. As large language models increasingly mediate the production of knowledge, they risk conflating eloquence with accuracy and replacing transparency with automation. The challenge ahead is therefore not only technical, improving factuality or retrieval, but institutional: ensuring that provenance, bias, and contestation remain visible in an age of synthetic knowledge. Future research should move beyond surface similarity to assess factual divergence, ideological asymmetry, and user trust. Comparing how readers evaluate credibility across human-edited and AI-generated sources may reveal whether Grokipedia’s expansion of text corresponds to an expansion, or erosion, of understanding.
5 Acknowledgement
This publication has emanated from research supported in part by grants from Taighde Eireann-Research Ireland under Grant numbers IRCLA/2022/3217. We acknowledge support from Workday, Inc.

### v4

@@ 3.6 Comparison within Topics
Because these articles cover a diverse set of subject areas, we additionally classified each article (see Methods for details). This allows us to compare similarity and bias shift not only across individual articles but also across broader thematic domains.
@@
Figure 7: Topic-level comparison of similarity and political bias shift between Wikipedia and Grokipedia articles. (a) Mean combined similarity scores between Wikipedia and Grokipedia articles within each topic, with standard-error bars. (b) Mean political bias shift (Grokipedia Wikipedia) for articles in each topic, indicating whether Grokipedia content tends to lean more left or right relative to Wikipedia. Error bars show standard errors within each topic.
@@ 4 Discussion and Conclusion
This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across 17,790 matched article pairs drawn from the 20,000 most-edited English Wikipedia titles. Despite their distinct production paradigms, community-driven versus AI-generated, the two encyclopedias exhibit remarkable alignment at the level of meaning and style. Semantic similarity averages around 0.86 and stylistic similarity around 0.88, indicating that Grokipedia reproduces much of Wikipedia’s linguistic and conceptual structure. However, the platforms diverge substantially in form and informational scaffolding. Grokipedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels and lexical diversity (greater syntactic complexity), but reference density. This pattern suggests that Grokipedia’s generation process elaborates on existing material, expanding text length and rhetorical flow, rather than producing substantively new or more rigorously sourced knowledge.
@@
The similarity distributions reveal additional nuance. The bimodal pattern across measures indicates that while some Grokipedia pages closely mirror Wikipedia’s organisation and tone, others diverge sharply. This heterogeneity suggests a hybrid authorship logic: part replication, part improvisation. In some cases, Grok appears to reproduce Wikipedia’s structure with high fidelity; in others, it diverges, generating content that is stylistically and semantically distinct.
Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [apnews2025, businessinsider2025, yasseri2025conversation]. Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia—often copied or paraphrased—while selectively emphasizing Musk’s personal and political narratives [rogers2025wired, oremus2025, hartwelle2025, wong2025]. Our quantitative results support this view: Grokipedia articles fall into two distinct clusters, with some heavily reliant on Wikipedia pages and others exhibiting systematic divergence.
@@
Complementing these textual and stylistic findings, our reference-level analysis (Figure 6) reveals that the magnitude and direction of political bias shifts between Grokipedia and Wikipedia articles are strongly related to article similarity. When Grokipedia reproduces Wikipedia closely (high similarity), the resulting entries show little or no bias shift. However, low-similarity articles display both greater variance and a rightward tendency in bias shift.
@@
In effect, Grokipedia trades collective accountability for computational authority. Even though some articles do diverge from Wikipedia, they still exhibit a left-leaning bias within (albeit less than their Wikipedia counterparts). However, the fact that this rightward shift is more pronounced in specific contested articles might point to the fact that this is not the result of the LLM model generating more neutral sources but a systematic decision by the creators. This cannot be examined as the prompts and the parameters used by the model is not publicly available.
Together, these findings point to a deeper epistemological contrast. Wikipedia’s openness renders bias visible and contestable through edits, disputes, and deliberation [yasseri2025conversation]. Grokipedia replaces this process with opaque, automated authorship, embedding potential bias within model behavior rather than exposing it to scrutiny. Despite its stated corrective aim, Grokipedia functions less as an epistemic alternative than as an AI-mediated reconfiguration of Wikipedia, trading collective accountability for computational authority.
Several limitations qualify these findings. First, the dataset focuses on Wikipedia’s 20,000 most-edited pages, of which 17,790 have counterparts on Grokipedia. This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [yasseri2012dynamics]. Second, the similarity metrics employed, lexical, semantic, structural, and stylistic, quantify form and textual alignment but not factual accuracy or ideological framing. The presence of hallucinated claims, selective omissions, or subtle rhetorical shifts remains outside the scope of automated comparison. Third, both platforms are dynamic: Wikipedia continuously evolves, while Grokipedia’s generative parameters may change with future model retraining. Finally, Grokipedia’s underlying data sources and editorial interventions are opaque, preventing full provenance auditing or causal inference about model bias.
@@
These findings extend beyond a single platform. As large language models increasingly mediate the production of knowledge, they risk conflating eloquence with accuracy and replacing transparency with automation. The challenge ahead is therefore not only technical, improving factuality or retrieval, but institutional: ensuring that provenance, bias, and contestation remain visible in an age of synthetic knowledge.
As AI-generated reference content may itself be used to train future language models, these patterns have implications beyond a single platform. Without transparency, accountability, and sustained human oversight, such derivative systems risk propagating existing biases into the epistemic foundations of future AI systems.
Future research should move beyond surface similarity to assess factual divergence, ideological asymmetry, and user trust. Comparing how readers evaluate credibility across human-edited and AI-generated sources may reveal whether Grokipedia’s expansion of text corresponds to an expansion, or erosion, of understanding.
Acknowledgement
This publication has emanated from research supported in part by grants from Taighde Eireann-Research Ireland under Grant numbers IRCLA/2022/3217 and 18/CRT/6049. TY acknowledge support from Workday, Inc.
Data Availability
Data used in the study are available at 10.5281/zenodo.18433470.